A community hall has an area of 396 metres squared. A 12 metre by 16 metre rug is placed on the floor so that the width of the floor showing on the four sides is uniform. What is the distance between the rug and the wall ?
Answer by checkley71(8403)

396=(12+2X)(16+2X)
X BEING THE SPACE BETWEEN THE RUG & THE WALL.
IT HELPS TO DRAW A DIAGRAM TO CAPTURE THE PROGRAM PICTORALLY.
396=192+32X+24X+4X^2
4X^2+56X+192-396=0
4X^2+56X-204=0
4(X^2+14X-51)=0
4(X-3)(X+17)=0
X-3=0
X=3 METERS IS THE SPACE BETWEEN THE RUG & THE WALL.